### v3

# Efficient Frame Extraction: A Novel Approach Through Frame Similarity and Surgical Tool Tracking for Video Segmentation

Abstract
The interest in leveraging AI for surgical procedures to automate analysis has witnessed a significant surge in recent years. One of the primary tools for recording surgical procedures and conducting subsequent analyses, such as performance assessment, is through videos. However, these operating room (OR) videos tend to be notably lengthy compared to other fields, spanning from thirty minutes to more than one hour, which poses a challenge for AI models to effectively learn from them. Despite this challenge, the foreseeable increase in the volume of such videos in the near future necessitates the development and implementation of innovative techniques to tackle this issue effectively. In this article, we propose a novel technique called Kinematics Adaptive Frame Recognition (KAFR) that can efficiently eliminate redundant frames to reduce dataset size and computation time while retaining useful frames to improve accuracy. Specifically, we compute the similarity between consecutive frames by tracking the movement of surgical tools. Our approach follows these steps: $i)$ Tracking phase: a YOLOv8 model is utilized to detect tools presented in the scene, $ii)$ Similarity phase: Similarities between consecutive frames are computed by estimating variation in the spatial positions and velocities of the tools, $iii$) Classification phase: A X3D CNN is trained to classify segmentation. We evaluate the effectiveness of our approach by analyzing [..111removed: a dataset obtained through a retrospective review of cases spanning from ] datasets obtained through retrospective reviews of cases at two referral centers. The Gastrojejunostomy (GJ) dataset covers procedures performed between 2017 to 2021 [..222removed: at two referral institutes. Within this ] , while the Pancreaticoduodenectomy (PJ) dataset spans from 2011 to 2022 at the same centers. In the GJ dataset, each robotic [..333removed: Gastrojejunostomy (GJ ) video was ] GJ video is segmented into six distinct phases. By adaptively selecting relevant frames, we achieve a tenfold reduction in the number of frames while increasing accuracy by three percent. Our approach is also evaluated on the PJ dataset, demonstrating its efficacy. In addition, we also compare our approach with [..444removed: a ] the state-of-the-art [..555removed: approach to demonstrate its superiority ] approaches to highlight its competitive in terms of performance and efficiency. Although we examined our approach on the GJ [..666removed: dataset ] and PJ datasets for phase segmentation, this could also be applied to broader, more general datasets. Furthermore, KAFR can be used as a supplement to existing approaches, enhancing their performance by reducing redundant data while retaining key information, making it a valuable addition to other AI models. Our source code is ready for replication at https://github.com/leonlha/Kinematics-AFR.
This paper includes surgical images that may be distressing to some readers.
Index Terms:
Adaptive Frame Recognition, Surgical Phase Segmentation, Tool Tracking, Convolutional Neural Networks, Deep Learning
[…]
=-21pt
I Introduction
The growing application of AI in video analysis is primarily inspired by developments in modern Graphics Processing Units (GPUs) [phong2022pso]. Traditionally, video analysis presented challenges due to their high dimension, temporal dynamics and visual variation. However, the emergence of AI has enabled researchers to develop increasingly intricate and data-driven approach for interpreting video. This has led to considerable performance increases for a variety of applications, for example, object detection [liu2023yolo, liang2015recurrent, felzenszwalb2009object], action recognition [sun2024k, phong2023pattern] and video segmentation [garg2024human, grammatikopoulou2024spatio]. AI models can effectively learn complex patterns and temporal correlations directly from raw video data, revolutionizing the field and opening up new opportunities for applications in surveillance [li2020abnormal], autonomous vehicles [gao2024human] and healthcare [IJMESD400].
In the field of surgery, there is an increasing trend towards leveraging advancements in AI, particularly for automating the analysis of surgical processes extracted from video recordings  [demir2023deep]. However, the duration of surgery videos might spans from thirty minutes to several hours, posing a significant challenge for AI models [..777removed: . Initially, statistical models ] , particularly in processing and analyzing lengthy sequential data effectively.
Early approaches to handling this challenge leveraged statistical models designed for sequential data, such as conditional random field [quellec2014real] , and hidden markov models [padoy2008line] [..888removed: were utilized ]. However, due to the complex temporal relations among frames, these techniques exhibit limited representation capacities with predefined dependencies [jin2017sv].
Nonetheless, as deep learning technology evolved and gained prominence in the research community, a plethora of novel techniques have emerged, achieving impressive outcomes. When dealing with sequential data, models such as Gated Recurrent Unit GRU [valipour2017recurrent], integration of Long Shot-Term Memory (LSTM) and Convolutional Neural Netwoks (CNN) [..999removed: can be utilized [jalal2021deep]. Beyond-Short-Snippets [yue2015beyond] aimed to extract feature maps from 2-D ConvNets and integrate them with LSTMs for video predictions. The Gated Recurrent Unit (GRU) was introduced with the aim of efficiently handling temporal data in the context of online segmentation [valipour2017recurrent]. Alternative approaches have also utilized ]  [jalal2021deep], and bi-directional LSTMs [..101010removed: , which consist of two distinct LSTMs, to explore both forward and backward temporal information ] can be utilized [he2021db].
[..111111removed: Though, LSTM models ] LSTM models, however, encounter the vanishing gradient problem when faced with long sequences of data. Alternatively, 3-D CNN models [al2024development, hegde2024automated] are favored for processing sequential data. This preference stems from their capacity to capture both spatial and temporal information simultaneously. A few notable examples illustrate this approach: Fractioned Adjacent Spatial and Temporal (FAST) 3-D convolutions are used to extract spatio-temporal patterns from 3-D videos, expanding on 2-D image convolutions [stergiou2019spatio]. CNN models have effectively distinguished between normal and Alzheimer’s disease patients utilizing an auto-encoder architecture and 3-D CNNs [hosseini2016alzheimer]. A fusion network that integrates the 3-D convolutional posture stream with the 2-D convolutional stream enhances the accuracy of human action recognition [huang2018human].
However, 3-D CNN can become bulky and challenging to train when using a large number of frames.
Vision transformers [nguyen2023video, han2023flatten, liu2021swin, arnab2021vivit] can handle longer sequences than CNN-LSTM models, but they require a large amount of data to train efficiently, which is often scarce and expensive in surgery.
[..121212removed: Hence, a ] Since surgical videos are lengthy and complex, increasingly advanced techniques are needed to manage the intricate temporal dependencies. To address this issue, our approach involves fine-tuning by identifying key frames and removing non-essential frames, thereby reducing the number of frames for training. Focusing on these key frames might help to train the network more efficiently by lowering computational costs, and speeding up processing. Specifically, in this approach, (1) background noise is filtered out while selectively preserving crucial surgical tools. (2) Adaptive 1 and Adaptive 2 methods are proposed to select most important frames. (3) Distinct X3D CNN models are used to train on the selected frames and corresponding optical flow data before ensembling.
The primary contributions of this paper include:
  1. 1.
    Introducing the use of kinematics data (velocity and acceleration) for adaptively selecting key frames in video-based phase segmentation in surgery, an crucial approach in this field.
  2. 2.
    Presenting a novel adaptive method that achieves state-of-the-art results.
  3. 3.
    Developing and annotating a new and comprehensive GJ dataset
  4. 4.
    Conducting extensive experiments on GJ dataset and PJ dataset to validate the effectiveness of the proposed methods.
  5. 5.
    The source code for our proposed methods is available for reproduction on GitHub at github.com
The rest of the article is organized as follows: In Section II, we discuss relevant approaches in applying Deep Learning in Surgery and selecting key frames. In addition, the proposed methods including Adaptive 1 and Adaptive 2 are introduced in Section III. The results of experiments and the extension of the experiments are presented in Sections V, and VI. Finally, we conclude our work in Section VIII.
II Related Works
II-A Surgery Datasets
Despite the growing interest in studying surgical skill performance, the number of public datasets remains limited. Unlike datasets in other fields, e.g., computer science, surgical datasets often contain sensitive patient information that must be kept confidential. Nevertheless, a few open datasets are available for research (See Table II-A).
TABLE I: List of surgery datasets.
{tblr}
row1 = Silver, column2 = c, column3 = c, Name & Videos Phases Description
Cholec80 [twinanda2016endonet] 80 7 Cholecystectomy procedures
from INRIA
Cholec51 [twinanda2016endonet] 51 7 A subset of Cholec80
CATARACTS
 [ALHAJJ201924] 50 19 Cataract procedures from
INRIA
M2cai16 [cadene2016m2cai] 41 8 Minimally invasive surgical
procedures from the
M2CAI 2016 Challenge
JIGSAWS [gao2014jhu] 17 4 Simulation procedures from
JHU
PJ [al2024development] 100 6 Pancreaticojejunostomy
GJ 42 6 Gastrojejunostomy
(this work)
One of the earliest surgical datasets, Cholec80, contains 80 videos of cholecystectomy surgeries performed by 13 surgeons. The dataset is divided into a training set (40 videos) and a testing set (40 videos), with the videos segmented into 7 phases. A subset of Cholec80, the Cholec51 dataset, includes 51 videos. However, our datasets differ as they focus on the procedures of Gastrojejunostomy and Pancreaticoduodenectomy, rather than cholecystectomy.
Additional datasets include the M2CAI16 Workflow Challenge dataset, which consists of 41 laparoscopic cholecystectomy videos (27 for training and 14 for testing), segmented into 8 phases. The CATARACTS dataset, with 50 videos, features 19 phase categories. We collected a similar number of videos (GJ with 42 videos and PJ with 100 videos).
The JHU-ISI Gesture and Skill Assessment Working Set (JIGSAWS) contains 39 videos performed by 8 users, with each video recorded approximately five times. While JIGSAWS is frequently used in surgical skill assessment research, it focuses on simulation tasks, whereas our dataset was recorded in an operating room during actual surgeries.
The PJ dataset, which we collected previously, comprises 100 videos segmented into 6 phases of Pancreaticoduodenectomy procedures. We use this dataset to provide a comparative baseline for our current research.
II-B Key Frame Recognition
As mentioned in the previous section, surgical videos tend to be longer than those in other fields, which poses a challenge for effectively using deep learning methods. A non-trivial approach entails significantly reducing the number of frames, for example, through Uniform Frame Sampling (UFS) which skips frames at regular intervals. Early attempts for key frame extraction relied on shot boundary-based algorithms [boreczky1996comparison]. Essentially, this technique utilizes the first or middle frame of each shot as the key frame after shot boundary detection. Although shot boundary-based approaches are simple to apply and generalize, the extracted key frame cannot fully capture the visual content. Another example used a Long Short-Term Memory (LSTM) network linked to the output of the underlying CNN [donahue2015long]. However, this framework obtains sixteen sample frames evenly split with an eight-frame stride from the full-length videos as the video representation. Although straightforward, this technique disregards the significance of consecutive frames, implying that certain frames hold significantly more importance than others.
One solution involves computing the similarity between frames and discarding those that exceed a certain threshold. However, strategies that are effective in other fields may not be applicable in surgery. For example, tools for minimally invasive surgery, such as scissors and forceps, often occupy a small portion of the screen and are overshadowed by background noise. It is worth noting that human organs are constantly in motion, presenting significant challenges that render approaches like background subtraction, optical flow, and even depth estimation, as well as non-deep learning methods such as Mean Square Error (MSE) and SSIM (which use all pixels of the image), less efficient [savran2023novel, yang2021fast, zhong2020key, cucchiara2003detecting].
[..131313removed: Our approach is based on a fine-tune process of filtering out background noise while selectively preserving those that play a crucial role in identifying key frames. By prioritizing the retention of only the most relevant elements, we aim to improve the efficiency of our method. This strategic refinement not only reduces the computational demands placed on our system but also enhances the precision and accuracy of our results. In essence, our strategy optimizes the balance between computational efficiency and analytical reliability, hence increasing the overall technical efficacy of our approach. ]
[..141414removed: The primary contributions of this paper include: ] [..151515removed: Introducing the use of kinematics data (velocity and acceleration)for adaptively selecting key frames in video-based phasesegmentation in surgery, an crucial approach in this field. ] [..161616removed: Presenting a novel adaptive method that achieves state-of-the-art results. ] [..171717removed: Developing and annotating a new and comprehensive GJ dataset ] [..181818removed: Conducting extensive experiments on GJ dataset to validate the effectiveness of the proposed methods. ] [..191919removed: The source code for our proposed methods is available for reproduction on GitHub at github. com ]
III Proposed Methods
III-A Architecture
Figure 1: Overview of the KAFR architecture. It involves three phases: (1) Object Tracking – Detects and tracks surgical tools. (2) KAFR – Computes centroids and identifies critical frames using tool movement. (3) Phase Segmentation – Classifies frames into segments with two X3D CNN models in a two-stream convolutional networks configuration.
Figure 1 presents an overview of our proposed architecture for KAFR. This workflow is divided into three distinct steps: Object Tracking, KAFR, and Phase Segmentation. The first phase – Object Tracking – plays an important role because the module traces the locations of surgical tools, which is critical for computing KAFR, especially when the tools are surrounded by noisy environment. In this phase, we extract the bounding box, class ID (with a confidence score greater than $0.5$), and frame number from the input video for surgical instruments. In the second phase, we compute the centroid of each tool, enabling us to track its movement. Please note that only frames from the training set were chosen. With the tools detected as polygon shapes, the centroids were determined using GeoPandas library in PyTorch. The distance and velocity are utilized as indicators to detect whether the surgeon’s maneuvers are essential. Subsequently, critical frames are identified. In the third phase, we employ two distinct X3D CNN models to classify frames into six different segments.
III-B Kinematics Adaptive Frame Recognition
III-B1 Adaptive 1
In a video sequence of frames $\{x_{1},x_{2},...,x_{n}\}$, a threshold $d$ determines the distance measure $D(x_{i},x_{j})$ used to classify a pair of frame $(x_{i},x_{j})$ as key frame or similar frame. Key frame pairs are denoted as $P_{\text{key}}$. Frames $x_{i}$ and $x_{j}$ are considered key frames if their distance is less than or equal to $d$. The frames between the two key frames are similar $P_{\text{similar}}$. The search for key frame pairs start from first frame until the last frame is encountered. Thus, we define a set of all key (K) pairs as follows:
$K(d)=\{(i,j)\mid D(x_{i},x_{j})\leq d,(i,j)\in P_{key}\}{\color[rgb]{0,0,1}\sf,}$ (1)
Assuming $S$ is a set of points of a frame and $s$ is a subset of $S$, the distance $D(x_{i},x_{j})$ are defined as
$D(x_{i},x_{j})=f\left(\sum\limits_{s\in S}\sum_{k=i+1}^{j}\left|\left|s(x_{i})-s(x_{k})\right|\right|\right),$ (2)
[…]
$f(z_{d})=\frac{1}{\left(z_{d}+\epsilon\right)^{\beta_{d}}},$ (3)
for constant $\beta_{d}>0$ and $\epsilon$ is introduced to avoid division by zero.
Figure 2: Illustration of Adaptive Frame Recognition. The subset $s$ consists of two tools, $M$ and $N$ (represented as circle and diamond shapes), [..232323removed: emerged ] emerging from a noisy environment (which might include organs of the abdominal cavity, other surgical tools, and noise, represented by other shapes). The tools are tracked across frames. The frame index $k$ is incremented one at a time until a designated threshold $d$ is [..242424removed: achieved ] reached.
III-C [..252525removed: Alternative KAFR ]
III-B1 Adaptive 2
One alternative to the equation (1) is using variation of velocity rather than distance
$K(d)=\{(i,j)\mid V(x_{i},x_{j})\leq d,(i,j)\in P_{key}\}{\color[rgb]{0,0,1}\sf,}$ (4)
Assuming $S$ is a set of points of a frame and $s$ is a subset $S$. The variation of velocity $V(x_{i},x_{j})$ is denoted as
$V(x_{i},x_{j})=f\left(\sum\limits_{\begin{subarray}{c}s\in S\end{subarray}}\sum_{k=i+1}^{j}\left|V_{\text{s}}(x_{i})-V_{\text{s}}(x_{k})\right|\right){\color[rgb]{0,0,1}\sf,}$ (5)
where $V_{\text{s}}(x_{k})$ is the velocity of point $x_{k}$ in the subset s
In Adaptive 2, we assume that
$f(z_{v})=\frac{1}{\left(z_{v}+\epsilon\right)^{\beta_{v}}}.$ (6)
for constant $\beta_{v}>0$ and $\epsilon$ is introduced to avoid division by zero.
IV [..272727removed: Experiments ]
III-A [..292929removed: Benchmark Dataset ] Object Detection
[..313131removed: Pancreatic cancer is projected to become the second leading cause of cancer-related deaths in the USA within the next two to three decades [mizrahi2020pancreatic]. Robotic Pancreaticoduodenectomy (also known as the Whipple procedure) surgery, which is a key procedure for treating pancreatic cancer, involves two main phases: Resection and Reconstruction. The Reconstruction phase includes three anastomoses: Hepaticojejunostomy, Pancreaticojejunostomy, and Gastrojejunostomy. ]
[..323232removed: A retrospective case review spanning from 2017 to 2021 was conducted at two prominent referral centers, UT Southwestern and University of Pittsburgh Medical Center. During this review, each robotic GJ video was segmented into six specific tasks: 1.1 Stay suture, 1.2 Inner running suture, 1.3 Enterotomy, 2.2 Inner running suture, 3.1 Inner Layer of Connell, and 4.1 Outer Layer of Connell, along with idle time. These videos were annotated by experts using a custom annotation framework. ]
[..333333removed: Of the 42 videos included, 30 were used for training and 12 for validation, with all frames extracted (six frames/second) and annotated, resulting in a dataset of 46,903 images for training and 22,647 images for validation. ]
III-B [..343434removed: Distribution of Tools ]
[..363636removed: Figure 7 displays the distribution of the number of frames for each class ID corresponding to the sixteen objects listed in Table II. We used YOLOV8 [redmon2016you], as described in Section III-B, to extract and track surgical instruments. Each instrument is segmented into three parts, each assigned a unique ID: the shaft, wrist, and jaw, except for the “laptool”. It’s important to note that in our dataset, instrument number two (“instrument2”) is identical to instrument number three (“instrument3”) because they are both needle drivers; therefore, they are combined under the same object category. ] [..373737removed: hbtp! ] [..383838removed: Distribution of Class IDs. Total of sixteen objects are tracked. ]
III-B [..393939removed: Object Detection ]
[..414141removed: The research utilized the ] As mention earlier, the detection of surgical tools is crucial. Therefore, we propose using the YOLOv8 neural network, the latest advancement in the “You Only Look Once” (YOLO) versions and is well-known for its real-time object detection capabilities, which outperform those of fast and faster Region CNN (RCNN) [girshick2014rich] for tool detection (See Figure 3). While YOLO comes pre-trained on general datasets like COCO [lin2014microsoft] or ImageNet [deng2009imagenet], its direct application in surgical video analysis may be suboptimal. Surgical instruments and operating environments possess unique characteristics, necessitating fine-tuning of the model for optimal performance.
Figure 3: [..444444removed: LabelMe for surgical tools annotation ] Tools auto-detection using Yolo.
We used a dataset comprises of 28 surgical videos and employed the LabelMe [russell2008labelme] tool to generate annotations, yielding 917 annotated images [..454545removed: . ] (See Figure 4). To the best of our knowledge, there were no other datasets containing the same set of surgical tools as those in the GJ and PJ datasets. Therefore, we used a portion of the GJ videos for tool segmentation. Although the recognition of individual surgical tools is distinct from phase segmentation, these two tasks are fundamentally different and should not affect each other’s performance. In addition, LabelMe automatically labels any part of an image that is not explicitly masked by the user as “background”. This classification is crucial for training detection models because it trains them to distinguish between objects of interest and irrelevant areas.
Figure 4: LabelMe for surgical tools annotation.
YOLO framework splits the input image into a grid of $S\times S$ cells. Each cell generates candidate bounding boxes and confidence levels for instrument or tools presence. The confidence metric Conf(Tool) is determined by the binary indicator Pr(Tool) and is defined as: [..464646removed: tbp! ] [..474747removed: Tools auto-detection using Yolo. ]
[…]
where Pr(Tool) denotes the presence of a tool (0 for absent, 1 for present). The Intersection Over Union (IOU) between the predicted box and the true box is calculated as:
$\text{IOU}=\frac{\text{area}(box_{\text{Truth}}\cap box_{\text{Pred}})}{\text{area}(box_{\text{Truth}}\cup box_{\text{Pred}})}$ (8)
III-B Surgical Phase Segmentation
III-B1 X3D Convolutional Neural Networks
Figure 5: Render X3D model architecture. (A) The X3D network extends a 2-D ResNet (res) one axis at a time across various axes, including temporal duration $X_{t}$, frame rate $X_{T}$, spatial resolution $X_{s}$, width $X_{w}$, bottleneck width $X_{b}$, and depth $X_{d}$. (B) Details of the 2-D ResNet demonstrate residual learning (Adapted from [feichtenhofer2020x3d] and [he2016deep]).
Surgical phase segmentation is an important aspect of computer-assisted surgery and medical image analysis, involving the automated partitioning of surgical procedures into distinct phases or stages using video or image data. The primary goal of surgical phase segmentation is to comprehend the temporal progression of surgeries, facilitating various clinical applications such as workflow analysis and skill assessment [tran2017phase].
To identify these phases, we utilized the X3D architecture, a highly efficient 3-D CNN designed for activity recognition and video classification (refer to Figure 5). X3D is an extension of the 2-D ResNet model originally designed for image classification, using residual learning to preserve essential information as it passes through layers.
III-B2 Two-stream Convolutional Networks
We integrated optical flow (OF) into our approach for video action recognition because of its critical role in capturing motion dynamics and temporal changes. Optical flow estimation involves computing angle and magnitude shifts for each pixel across consecutive video frames. The Farneback [farneback2003two] algorithm is a common technique for dense optical flow computation, that utilizes polynomial expansion and iterative field estimation. By solving a quadratic polynomial equation between two consecutive frames $I(x,y,t)$ and $I(x+\Delta x,y+\Delta y,t+\Delta t)$, optical flow $(u,v)$ at each pixel is estimated to obtain displacement fields.
In our work, we implemented the Farneback optical flow estimation using the calcOpticalFlowFarneback function from OpenCV. We set the number of pyramid levels to $3$, the pyramid scale to $0.5$, the window size to $15$, the number of iterations to $3$, the pixel neighborhood size to $5$, and a standard deviation of $1.2$. The process includes converting frames to gray-scale, using the calcOpticalFlowFarneback function to compute flow vectors, and saving the magnitude as an image. The estimated optical flow may be utilized as input to our X3D CNN model.
Back to the Figure 1, in the Phase Segmentation module, we trained two distinct X3D models separately. In the last layer of both models, instead of outputting classifications, we modified the layer to output probabilities. We then ensembled these probabilities to make classifications.
IV Experiments
IV-A Benchmark Dataset
Pancreatic cancer is projected to become the second leading cause of cancer-related deaths in the USA within the next two to three decades [mizrahi2020pancreatic]. Robotic Pancreaticoduodenectomy (also known as the Whipple procedure) surgery, which is a key procedure for treating pancreatic cancer, involves two main phases: Resection and Reconstruction. The Reconstruction phase includes three anastomoses: Hepaticojejunostomy, Pancreaticojejunostomy, and Gastrojejunostomy.
IV-A1 GJ Dataset
A retrospective case review spanning from 2017 to 2021 was conducted at two prominent referral centers, UT Southwestern and University of Pittsburgh Medical Center. During this review, each robotic GJ video was segmented into six specific tasks: 1.1 Stay suture, 1.2 Inner running suture, 1.3 Enterotomy, 2.2 Inner running suture, 3.1 Inner Layer of Connell, and 4.1 Outer Layer of Connell, along with idle time. These videos were annotated by experts using a custom annotation framework (See Figure 6.
[ [..484848removed: tbhp! ] hbtp!]
[..494949removed: Render X3D model architecture. (A) The X3D network extends a 2-D ResNet (res) one axis at a time across various axes, including temporal duration $X_{t}$, frame rate $X_{T}$, spatial resolution $X_{s}$, width $X_{w}$, bottleneck width $X_{b}$, and depth $X_{d}$. (B) Details of the 2-D ResNet demonstrate residual learning (Adapted from [feichtenhofer2020x3d] and [he2016deep]). ] Refer to caption
Figure 6: Gastrojejunostomy Steps.
IV-B [..505050removed: Architecture ]
[..525252removed: Figure 1 presents an overview of our proposed architecture for KAFR. This workflow is divided into three distinct steps: Object Tracking, KAFR, and Phase Segmentation. In the first phase, we extract the bounding box, class ID (with a confidence score greater than $0.5$) ] Of the 42 videos included, 30 were used for training and 12 for validation, with all frames extracted (six frames/second) and annotated, resulting in a dataset of 46, [..535353removed: and frame number from the input video for surgical instruments. In the second phase, we compute the location or centroid of a tool, enabling us to track its movement. The distance and velocity are utilized as indicators to detect whether the surgeon’s maneuvers are essential. Subsequently, critical frames are identified. Afterwards, we employ X3D CNN to classify frames into six different phases. ] 903 images for training and 22,647 images for validation.
IV-A1 PJ Dataset
A retrospective review was conducted on recording videos collected from 2011 to 2022 across two quaternary referral centers: UT Southwestern Medical Center and the University of Pittsburgh Medical Center [al2024development]. Out of the 100 videos reviewed, 60 were utilized for model training, 10 for hyperparameter optimization, and 30 for performance testing. Frames were extracted at 6 frames/second and annotated. Each video was divided into tasks, with start and end times noted for each. Six tasks were annotated per video: 1.1 Anterior Mattress Sutures, 1.2 Tying Mattress Sutures, 2.1 Enterotomy, 3.1 Posterior Duct Sutures, 3.2 Anterior Duct Sutures, and 4.1 Anterior Buttress Sutures.
IV-B Distribution of Tools
Figure 7 displays the distribution of the number of frames for each class ID corresponding to the sixteen objects listed in Table II. We used YOLOV8 [redmon2016you], as described in Section III-B, to extract and track surgical instruments. Each instrument is segmented into three parts, each assigned a unique ID: the shaft, wrist, and jaw, except for the “laptool”. It’s important to note that in our dataset, instrument number two (“instrument2”) is identical to instrument number three (“instrument3”) because they are both needle drivers; therefore, they are combined under the same object category.
Figure 7: [..575757removed: Proposed Architecture ] Distribution of [..585858removed: KAFR ] Class IDs. Total of sixteen objects are tracked.
TABLE II: Tool name. The name of tools along with its parts and ids.
{tblr}
row1 = Silver, hlines, vlines, ID Instrument Part Tool Name
0 instrument3part1 Needle Driver
1 instrument3part2 Needle Driver
2 instrument3part3 Needle Driver
3 instrument2part3 Needle Driver
4 instrument2part2 Needle Driver
5 instrument2part1 Needle Driver
6 laptool Irrigator
7 instrument4part1 Forcep
8 instrument1part1 Grasper
9 instrument1part2 Grasper
10 instrument1part3 Grasper
11 instrument4part2 Forcep
12 instrument5part3 Monopolar Curved Scissors
13 instrument5part2 Monopolar Curved Scissors
14 instrument5part1 Monopolar Curved Scissors
15 instrument4part3 Forcep
IV-C Hyper-parameter
[..595959removed: The experiments in this research ] Our proposed methods were developed using a PyTorch framework and implemented on a Windows (R) 11 Enterprise system equipped with an AMD RyzenTM ThreadripperTM PRO 3995WX 64-cores CPU @ 2.70 GHz. We trained our model with two Nvidia RTX A6000 GPUs.
The experiments were performed based on the configurations specified in Tables III. The hyperparameters such as image size, batch size, and clip length were optimized to make full use of GPU memory. Please note that the threshold values for KAFR are determined based on the desired number of extracted frames. Instead of selecting a value between 0 and 1 as in the Equation 1, we choose a number of frames by a percentage of total number of training samples, and adjust the threshold accordingly.
TABLE III: Hyperparameters.
Hyperparameter Value
X3D CNN
Image size $300\times 300$
Optimizer AdamW
Schedule StepLR
Learning rate 0.001
Decay rate 0.7
Batch size 64
Epochs 100
Early stop 10
Clip length 16
Number frame per step 250
KAFR
Threshold Flexible
IV-D Evaluation Metrics
We utilize traditional evaluation metrics such as accuracy and F1 score to analyse the performance of our models.
  1. 1.
    Accuracy: it computes the percentage of correct predictions over all predictions made on the test data; that is, it represents the proportion of predictions that were correctly classified relative to the total number of predictions.
    Accuracy=Number of correct predictionsTotal number of $\text{Accuracy}=\frac{\text{Number of correct predictions}}{\text{Total numberof predictions}}$ (9)
  2. 2.
    F1 score: it establishes a uniform metric for evaluating the precision and recall by taking into account both false positives and false negatives.
[…]
V Classification Results
V-A Adaptive 1
In this experiment, we initially focused on tracking velocity using a single tool—specifically, the right-hand tool, as most people are comfortable with their right hand. However, this method frequently fell short of surpassing the baseline accuracy of $0.749$ (During our experiments, we identified the optimal parameter that yielded the most favorable results and this “baseline” represents a best finding or achievement derived from our research efforts). The turning point occurred when we allocated 15% of our data to the task, resulting in a substantial improvement that exceeded the baseline almost $0.22\%$ (See Figure 8 and also Table IV). Encouraged by this accomplishment, we increased our efforts to refine our approach further.
Building upon these promising findings, we expanded our experiments to incorporate velocity tracking from two tools (left and right hand). The result was remarkable; with just a 10% data allocation, we achieved a $1.9$% increase in accuracy. Our observations indicated that the method was most effective with sufficient data reduction (20%, 15%, and 10% of total frames), suggesting that critical frames were being effectively captured. Though, significant reductions in data (5% and 1% of total frames) led to decreased accuracy, likely due to insufficient data for robust model training.
Subsequent studies involving additional tools yielding slight accuracy gains (30%, 20%, 10%, and 5%), and in some cases, resulting in reduced accuracy (50%, 15%, and 1% data allocations). Though, our efforts achieved in a better accuracy of $0.7684$ by utilizing four tools with a 10% frame allocation.
These outcomes indicate the important role of adaptive learning in optimizing model performance. Our Adaptive 1 method, which selects frames based on recognition of significant tool movements, not only improves accuracy but also notably reduces the dataset size. In the next section, we will compare our approach with a well-known method to highlight its superiority.
Figure 8: Performance Using Velocity.
TABLE IV: Effect of Velocity.
{tblr}
row1 = Silver, row2 = Silver, column3 = c, column4 = c, cell11 = c=2, cell13 = c=2, cell31 = r=7, cell101 = r=7, cell171 = r=7, vlines, hline1-3,10,17,24 = -, hline4-9,11-16,18-23 = 2-4, Performance
Data Model Accuracy F1 Score
One Object X3D CNN ($1\%$) 0.7133 0.6071
X3D CNN ($5\%$) 0.7388 0.6806
X3D CNN ($10\%$) 0.7481 0.6857
X3D CNN ($15\%$) 0.7519 0.6830
X3D CNN ($20\%$) 0.7495 0.6829
X3D CNN ($30\%$) 0.7475 0.6733
X3D CNN ($50\%$) 0.7430 0.6577
Two Objects X3D CNN ($1\%$) 0.7501 0.6044
X3D CNN ($5\%$) 0.7538 0.6529
X3D CNN ($10\%$) 0.7680 0.7129
X3D CNN ($15\%$) 0.7535 0.6781
X3D CNN ($20\%$) 0.7635 0.6812
X3D CNN ($30\%$) 0.7544 0.6810
X3D CNN ($50\%$) 0.7592 0.6769
Four Objects X3D CNN ($1\%$) 0.7458 0.6002
X3D CNN ($5\%$) 0.7575 0.6816
X3D CNN ($10\%$) 0.7684 0.6900
X3D CNN ($15\%$) 0.7666 0.6943
X3D CNN ($20\%$) 0.7683 0.6906
X3D CNN ($30\%$) 0.7550 0.6906
X3D CNN ($50\%$) 0.7578 0.6754
V-B Comparison
This experiment provides a comparative analysis between our proposed method, Kinematics Adaptive Frame Recognition (KAFR), and Mean Squared Error (MSE). MSE, a widely used metric in tasks such as detection of moving object and loss function measurement, evaluate changes over time using all pixel-wise information within an image. In contrast, KAFR focuses just on the tracked tools, disregarding other pixel data. This distinction highlights a key difference between the two methods: MSE considers the entirety of pixel information, while KAFR concentrates solely on the identified tools.
Figure 9 illustrates the performance of KAFR and MSE relative to a baseline, indicating whether their accuracy has improved or decreased. MSE occasionally outperforms the baseline with 20%, 15%, and 5% data allocations but exhibits notable defeats with 50%, 30%, 10%, and 1% data allocations, implying vulnerability to background noise interference that impacts its performance. It should be emphasized that in surgical videos, the background which typically consists of deformable organs and connective tissues is constantly moving, making it harder to extract useful information from tool movements. On the other hand, KAFR outperforms the baseline by approximately $1.94$%, demonstrating effectiveness in selectively tracking and utilizing tool data to identify key frames. In addition to comparing to the baseline, when both methods are put together, KAFR performs better than MSE.
Figure 9: KARF vs. MSE.
V-C Adaptive 2
An alternative method to detect tool movement is to examine speed of changes in velocity, called acceleration. These accelerations may be a sign of significant surgical maneuvers made by the surgeons. Looking at the Figure 10, a trend emerges similar to that of velocity, where in most cases, the accuracy outperforms the baseline. However, the accuracy reduced when the number of frames is decreased to 5% and 1%. Although the significance of accelerations in detecting crucial frames may not match that of velocity, we achieved better accuracy ($0.771$) using a 15% data allocation for accelerations.
Figure 10: Comparison of Velocity, Acceleration, and Baseline.
V-D Two-stream Convolutional Networks
[..606060removed: We integrated optical flow (OF) into our approach for video action recognition because of its critical role in capturing motion dynamics and temporal changes. Optical flow estimation involves computing angle and magnitude shifts for each pixel across consecutive video frames. The Farneback [farneback2003two] algorithm is a common technique for dense optical flow computation, that utilizes polynomial expansion and iterative field estimation. By solving a quadratic polynomial equation between two consecutive frames $I(x,y,t)$ and $I(x+\Delta x,y+\Delta y,t+\Delta t)$, optical flow $(u,v)$ at each pixel is estimated to obtain displacement fields. ]
[..616161removed: In our work, we implemented the Farneback optical flow estimation using the calcOpticalFlowFarneback function from OpenCV. We set the number of pyramid levels to $3$, the pyramid scale to $0.5$, the window size to $15$, the number of iterations to $3$, the pixel neighborhood size to $5$, and a standard deviation of $1.2$. The process includes converting frames to gray-scale, using the calcOpticalFlowFarneback function to compute flow vectors, and saving the magnitude as an image. The estimated optical flow may be utilized as input to our X3D CNN model. ]
Table V shows the results achieved using the two best candidates from the previous section, namely Acceleration with 2-Tool using 10% of the data and Velocity with 4-Tool using 15% of the data. The outputs of the two channels (Optical Flow and RGB) are ensembled. As we can see, the accuracy is further improved to $0.7814$.
TABLE V: Two-stream Convolutional Networks.
{tblr}
row1 = Silver, row2 = Silver, column3 = c, column5 = c, cell11 = r=2, cell12 = c=2c, cell14 = c=2c, cell32 = c, cell34 = c, cell42 = c, cell44 = c, cell52 = c, cell54 = c, vlines, hline1,3-6 = -, hline2 = 2-5, Type X3D CNN Two Objects
Acceleration ($15\%$) X3D CNN Four Objects
Velocity ($10\%$)
Accuracy F1 Score Accuracy F1 Score
Optical Flow 0.7291 0.6496 0.7175 0.6304
RGB 0.7438 0.6747 0.7532 0.6750
Ensemble 0.7814 0.7141 0.7666 0.6794
Figure 11: Phase Segmentation. Two distinct videos are sampled to show KAFR performance. The color codes are: ’1.1 Stay suture’ - ’Blue’, ’1.2 Inner running suture’ - ’Yellow’, ’1.3 Enterotomy’ - ’Green’, ’2.2 Inner running suture’ - ’Orange’, ’3.1 Inner Layer of Connell’ - ’Indigo’, ’4.1 Outer Layer of Connell’ - ’Violet’, ’Idle time’ - ’Gray’, respectively.
Figure 11 illustrates the results (X3D CNN 2Tool Acceleration (10%)) using two distinct videos. As shown in the figure, six phases are segmented along with idle time (in gray color). The frame indices for the original data (showing gaps between frames resulting from undersampling and oversampling techniques, as more samples are extracted for shorter phases and fewer samples for longer phases) and the filtered indices after using KAFR are displayed. The KAFR tends to skip frames at the beginning and end of a phase, e.g., the first frames in gray color and the end frames in pink color in the upper figure. It is also observed that in the Enterotomy phase of the upper figure, the gap for filtered data appears longer. This is because, in this specific instance, the surgeon performed the cutting more slowly compared to other tasks in distinct phases.
VI Extension
In this section, we apply our proposed methods, Adaptive 1 and Adaptive 2, to the PJ dataset (See Section IV-A1). For consistency, we maintained the same model architecture (X3D CNN) and optimized settings as detailed in previous research [al2024development]. With this setup, the best model achieved an accuracy of 0.8801, which we use as the baseline for comparison with our findings.
VII [..626262removed: Discussion ]
[..646464removed: Our ] Figure 12 presents a comparison of the results obtained by our proposed methods against the baseline. In this evaluation, we examined the effect of varying the training set size, using increments of $1\%$, $5\%$, $10\%$, $15\%$, $20\%$, $30\%$, and $50\%$ of the training data. Our study also explored the impact of tracking data from multiple tools on model accuracy. Specifically, we analyzed scenarios where tracking was applied using one, two, and four tools in the same manner as with the GJ dataset.
Our best results outperformed the baseline, demonstrating the effectiveness of our adaptive methods. The highest accuracy for Adaptive 1 reached 0.8909 (a $1.01\%$ increase) when using velocity data from one tool with 30% of the data. For Adaptive 2, the top accuracy was 0.8976 (a $1.75\%$ increase), achieved with acceleration data from two tools and 20% of the data. The accuracy is increased to 0.8982 with the ensembling method, employing Adaptive 2 with 2-Tool approach on 20% of the data.
Figure 12: Accuracy on PJ Dataset. (A) Adaptive 1 – Velocity. (B) Adaptive 2 – Acceleration.
VII Comparison with the state of the art
Previously, our group conducted research on the PJ dataset using the X3D CNN model. In this study, we build upon that work by introducing a new method, Kinematics Adaptive Frame Recognition, to selectively capture key frames rather than using all frames. This approach reduces the number of frames by fivefold and improves accuracy by approximately 2%.
TABLE VI: State-Of-The-Art Methods on Distinct Datasets.
[…]
Cholec80 TeCNO [czempiel2020tecno] 88.56
Endonet + HHMM [twinanda2016endonet] 81.70
LoViT [liu2025lovit] 92.40
ST-ERFNet [jin2020multi] 86.07
MTRCNet-CL [jin2020multi] 87.40
Cholec51 TeCNO [czempiel2020tecno] 87.34
CATARACTS SurgPLAN [luo2024surgplan] 83.10
M2F + GCN [holm2023dynamic] 76.86
M2cai16 LAST [tao2023last] 91.50
PJ X3D CNN [al2024development] 88.01
PJ KAFR 89.82
GJ KAFR 78.14
The TeCNO represents a Multi-Stage Temporal Convolutional Network (MS-TCN) that uses hierarchical prediction refinement approach for surgical phase recognition. It employs causal, dilated convolutions, to provide a large receptive field, enabling smooth, online predictions even during ambiguous transitions in surgical phases. In contrast to standard acausal MS-TCNs, the TeCNO restricts each prediction to rely only on current and past frames, enhancing real-time application by reducing dependent on future frames. Applying this method to the Cholec80 and Cholec51 datasets yielded accuracies of 88.56% and 87.34%, respectively. Although TeCNO arguably surpasses many state-of-the-art LSTM approaches, training ResNet50 frame-by-frame without temporal context may reduce effectiveness by loosing crucial frame-to-frame relationships. Notably, its accuracy is worse than that of our KAFR approach on the PJ dataset.
The Long Video Transformer (LoViT) takes another approach by focusing the extraction of temporal spatial features and modeling phase transitions. It utilizes a multiscale temporal aggregator that integrates L-Trans modules based on self-attention to capture short-term, fine-grained details, and a G-Informer module utilizing ProbSparse self-attention to examine long-term temporal dynamics. These features are fused through a multiscale temporal head, which employs phase transition-aware supervision for accurate classification. The model achieves an accuracy of 92.40% on the Cholec80 dataset, slightly outperforming our result on the PJ dataset (89.82%). However, it does not take into consideration the dynamics of surgical tool movement, such as velocity and acceleration, which are addressed in Adaptive 1 and Adaptive 2. Additionally, the source code was unavailable at the time of writing this manuscript, preventing a fair comparison by applying the LoViT model to our datasets.
In terms of key frames recognition, our proposed method differs from a recent paper [..656565removed: in the field ] [loukas2018keyframe], although both aim to detect key frames within a video. First, our approach [..666666removed: identifies significant frames based on the movement of a limited set of tools, including their velocity and acceleration. Our model ]uses YOLOv8 to detect the presence of these tools in the scene and track their locations across frames. In contrast, the other approach relies on color, motion, and texture and uses a less fine-grained method for motion detection, tracking all moving objects without focusing on specific tools. Second, our goal is to use key frames for phase segmentation or gesture detection, which requires more frames than the other approach, whose objective is merely to select representative frames for summarizing the video, requiring fewer frames. Additionally, we use the X3D CNN model to analyze the temporal information of 2-D frame sequences instead of the HMMAR model.
VIII Conclusion
In conclusion, the growing interest in utilizing AI to automate surgical phase segmentation analysis has prompted extensive research efforts. Surgical videos, essential for performance assessment and analysis, pose challenges to AI models due to their length. However, with the expected increase in video volume, there is an urgent need for innovative techniques to address these challenges effectively.
In this study, we introduced Kinematics Adaptive Frame Recognition (KAFR), a novel method aimed at reducing dataset size and computation time by removing redundant frames while preserving critical information to improve accuracy. The KAFR approach consists of three primary phases: i) Use a YOLOv8 model to detect surgical tools in the scene; ii) Compute similarities between consecutive frames by analyzing spatial positions and velocities of detected tools; iii) Train a X3D CNN to perform phase segmentation based on extracted frames.
We evaluated the efficacy of KAFR on [..676767removed: a GJ ] both the GJ and PJ dataset over a span of [..686868removed: 6â ] €“ [..696969removed: 8 ] 11 weeks. By adaptively selecting relevant frames, we achieved a tenfold reduction in number of [..707070removed: frame ] frames while improving accuracy by more than 3% on GJ dataset and fivefold reduction in frames with an approximately 2% accuracy improvement on the PJ dataset. Furthermore, we conducted a comparative analysis with the state-of-the-art method MSE, to demonstrate the [..717171removed: superior ] competitive performance and efficiency of our approach. Moreover, KAFR can be used to enhance existing methods, providing a valuable improvement to the performance of other AI models.
Acknowledgments
We acknowledge the funding from NIH/NIBIB R01: EB025247 that supported Drs. Nguyen, Garces Palacios and Sankaranarayanan.
IX Ethical approval
The institutional review board – the ethics committee of the UT Southwestern Medical Center approved the study design and the use of GJ videos. All methods were performed in accordance with the relevant guidelines and regulations.
X Data Availability
The data that support the findings of this study may be provided upon reasonable request.
[Uncaptioned image] Huu Phong Nguyen received a BSc in Physics from Vietnam National University, Hanoi, an MSc in Information Technology from Shinawatra University, Thailand, and a PhD degree from the University of Coimbra, Portugal. He was a member of the Adaptive Computation group at the Centre of Informatics and Systems at the University of Coimbra (CISUC). Currently, he is pursuing a PostDoc at the University of Texas Medical Center, USA. His research interests include medical imaging, image classification, action recognition, pattern recognition, and deep learning.
[Uncaptioned image] Shekhar Madhav Khairnar is currently working as a Data Scientist at UT Southwestern Medical Center, focusing on developing AI models, particularly in computer vision domain. He earned a Bachelor of Engineering (BE) in Electrical from the University of Pune, India, and a Master of Science (MS) in Industrial Engineering from the University of Texas at Arlington, where he received awards including the GTA/GRA and G.T. Stevens Endowed Scholarship. His research areas include computer vision, covering image classification, segmentation, object detection, and tracking, and broader fields such as machine learning, natural language processing (NLP), large language models (LLMs), and deep learning.
[Uncaptioned image] Sofia Garces Palacios is a Research Associate at the Artificial Intelligent and Medical Simulation (AIMS) Laboratory within the Department of Surgery at UT Southwestern Medical Center. Dr. Sofia holds a Doctor of Medicine degree from the University of Manizales in Colombia and is pursuing a career in Surgery. Her research interests are centered on surgical education, minimally invasive surgical techniques, and global surgery.
[Uncaptioned image] Amr Al-Abbas is a General Surgery Resident at UT Southwestern Medical Center and Parkland Hospital, with over five years of experience. He also served as a Postdoctoral Research Fellow at UT Southwestern, focusing on artificial intelligence for video-based assessment in robotic pancreatic surgery. Previously, Dr. Al Abbas completed postdoctoral research in surgical oncology at UPMC and the University of Pittsburgh, where he conducted clinical outcomes research. He has additional research experience in bariatric surgery at the American University of Beirut. His academic background includes work in cell and molecular biology and surgical oncology sub-internships at Duke University and the American University of Beirut. Dr. Amr Al Abbas is interested in surgical oncology robotic, surgery surgical, education video-based assessment, machine learning.
[Uncaptioned image] Melissa E. Hogg is the Director of HPB Surgery at Northshore University Hospital and Director of Advanced Robotic Training for the Grainger Center for Simulation and Innovation. She did her medical school and residency training at Northwestern University in Chicago, IL, and her fellowship and masters training in Medical Education and Clinical Research at the University of Pittsburgh, PA where she also had her first faculty appointment. She is faculty of the University of Chicago Pritzker School of Medicine and serves as Associate Program Director of the General Surgery Residency.
[Uncaptioned image] Amer H. Zureikat is chief of the Division of Gastrointestinal (GI) Surgical Oncology at UPMC Hillman Cancer Center, co-director of the UPMC Hillman Cancer Center Pancreatic Cancer Program, and associate professor of surgery at the University of Pittsburgh School of Medicine. Dr. Zureikat specializes in cancers and diseases of the pancreas, stomach, liver, and duodenum, and practices state-of-the-art robotic surgery.
[Uncaptioned image] Patricio Polanco is an Associate Professor of Surgical Oncology at UT Southwestern Medical Center, specializing in minimally invasive and robotic surgery for pancreatic and liver cancers. He directs the Robotic Surgery Training Program and co-directs the Pancreatic Cancer Program and Prevention Clinic. Dr. Polanco leads the Peritoneal Surface Malignancies and HIPEC Program at the Harold C. Simmons Comprehensive Cancer Center. Originally from Peru, he completed his medical education at Universidad de San Martin de Porres and further specialized in complex surgical oncology in the U.S. He is double board-certified and has published over 100 peer-reviewed articles, earning numerous research awards. Dr. Polanco is a recognized leader in robotic surgery training and has served in various leadership roles within national and international surgical associations. His contributions to global health include training surgeons and participating in medical missions across Latin America.
[Uncaptioned image] Herbert J. Zeh III is Professor and Chair of the Department of Surgery at UT Southwestern Medical Center, where he holds the Hall and Mary Lucile Shannon Distinguished Chair in Surgery. A leader in pancreatic diseases and cancer, Dr. Zeh has over 20 years of surgical experience. He earned his medical degree from the University of Pittsburgh and completed his residency and fellowship at The Johns Hopkins Hospital. Before joining UT Southwestern in 2018, he served as the Watson Family Professor of Surgery at the University of Pittsburgh and co-directed the UPMC Pancreatic Cancer Center. A prolific author with over 200 peer-reviewed publications, Dr. Zeh’s research focuses on novel treatments for pancreatic cancer and robotic surgery innovations. He is a member of numerous prestigious surgical organizations.
[Uncaptioned image] Ganesh Sankaranarayanan received the BE degree in electrical electronics engineering from Annamalai University in 1997, the MS degree in electrical engineering from the University of Texas at Arlington in 2002, and the PhD degree in electrical engineering from the University of Washington in 2007. He was a postdoctoral research associate from 2008-2010 and research assistant professor from 2010-2013 in the Department of Mechanical, Aerospace, and Nuclear Engineering. He was a research scientist at the Center for Modeling, Simulation, and Imaging in Medicine (CemSIM) at the Rensselaer Polytechnic Institute. He is currently a clinical associate professor in the Department of Surgery at the Texas A&M Health Science Center and assistant director of the Center for Evidence Based Simulation (CEBS) at the Baylor University Medical Center in Dallas, TX. His interests are in haptic device design, surgical robotics, surgical simulation, and human computer interaction. He is a member of the IEEE and IEEE Robotics and Automation Society.
[…]